# Towards Scalable Language-Image Pre-training for 3D Medical Imaging

### v2

@@ 4 Experiments
We apply HLIP to brain MRI, head CT, and chest CT. For brain MRI, we construct a new benchmark for zero-shot classification, using public datasets [46, 23, 47, 48, 49, 50, 51, 52]. For head CT and chest CT, we follow evaluation protocols from previous work [3, 6, 20]. Experiments on our uncurated brain MRI and head CT datasets underscore the importance of effective modeling and demonstrating superior performance of HLIP over current foundation models [32, 14, 20, 53]. Experiments on curated chest CT datasets [24, 3] further isolate the effectiveness of the proposed hierarchical attention mechanism. Finally, we conduct comprehensive ablation and analysis to investigate our designs.
@@ 4.1 Brain MRI
Pub-Brain-5. To the best of our knowledge, no publicly available benchmark exists for zero-shot transfer evaluation on brain MRI tasks [^2] BrainMD [35] is not publicly available due to an unexpected privacy policy.. To this end, we construct a benchmark, named Pub-Brain-5, based on existing publicly available brain MRI datasets. Pub-Brain-5 comprises 18,343 studies drawn from Open-BHB [46], the Stroke dataset [23], BraTS 2023 [47, 48, 49, 50], NYU-Mets [52], and UCSF-Mets [51]. It spans five classes: healthy (3,984), acute stroke (2,871), glioma (1,614), meningioma (1,141), and metastasis (8,733). For each study, the number of scans ranges from 1 to 14. We also construct a subset of Pub-Brain-5, namely Pub-Brain-5-GT, which contains lesion-containing slice annotations derived from the original segmentation ground truth. Pub-Brain-5-GT comprising 8,944 studies drawn from Open-BHB, the Stroke dataset, and BraTS 2023, covering the same five classes: healthy (3,984), acute stroke (2,372), glioma (1,350), meningioma (1,000), and metastasis (238). We evaluate three zero-shot tasks: (i) binary anomaly detection (i.e., distinguishing diseased from healthy studies); (ii) three-way tumor classification; and (iii) five-way disease classification.
Baselines. We first evaluate the zero-shot transferability of two biomedical CLIP models on our benchmark: BiomedCLIP [32], pre-trained on 15 million figure-caption pairs from PubMed, and ConceptCLIP [14], pre-trained on 23 million such pairs. Since both models require 2D inputs, we generate study-level predictions by applying either average pooling or max pooling to the outputs across all slices. We observe that max pooling consistently produces positive predictions and is therefore only suitable for binary classification; accordingly, we adopt average pooling for both models. To enable a fairer comparison between these two models and HLIP, we assess an upper-bound scenario on Pub-Brain-5-GT, where predictions for the two models are made only on lesion-containing slices, while HLIP still takes the raw study as input. In addition, we evaluate PRIMA [54], which represents the vqvae + hierarchical transformer architecture family [3, 54], as well as a vanilla ViT that encodes the entire study.
Implementation Details. We use the prompt "This brain MRI shows: {disease}." to perform zero-shot inference [2] with biomedical CLIP models. HLIP and the vanilla ViT are trained on BrainMRI220K as described in Section 3.2, while PRIMA [54] is re-trained on the same dataset using the training recipe provided in the original paper. These three models employ the prompt "This MRI study shows: {disease.}" during zero-shot transfer.
@@
Method Anomaly Detection
Tumor
Disease
@@
BiomedCLIP [32] 64.7 87.8 63.6 59.8 69.0 50.4 31.5
ConceptCLIP [14] 66.8 91.9 57.7 67.9 71.1 35.7 30.9
PRIMA [54] 78.8 89.3 70.8 64.7 75.9 42.8 31.6
@@
Results. Given both benchmarks are imbalanced, we report balanced accuracy (ACC) in Table 1. On Pub-Brain-5, HLIP demonstrates superior performance over all baseline models, outperforming the second-best baseline by 20.5% ACC in disease classification. On Pub-Brain-5-GT, we observe a substantial performance boost for two biomedical CLIP models [32, 14] when predictions are restricted to lesion-containing slices, underscoring the fairness of our evaluation for these models even on uncurated studies. Notably, in disease classification, ConceptCLIP [14] achieves 50.8% ACC, outperforming PRIMA [54] and the vanilla ViT trained on our large-scale domain-specific datasets, indicating that merely scaling up data is not a sufficient solution for language-image pre-training in 3D medical imaging. On the other hand, HLIP achieves the new SOTA 61.3% ACC, attributable to the effectiveness of the proposed hierarchical attention mechanism.
@@ 4.2 Head CT
Implementation Details. We apply HLIP to HeadCT240K, as described in Section 3.2. For linear-probe evaluation, we benchmark HLIP on two public datasets, CQ500 [21] and RSNA [22], exactly following the protocol of FM-HeadCT [20]. For zero-shot evaluation, we evaluate HLIP on the full CQ500 and RSNA datasets and compare it with the vanilla ViT trained on the same dataset [^3] Recent CLIP models for 3D medical imaging, such as Merlin [4] and M3D [5], are excluded from our baselines, as they perform no better than random guessing in zero-shot evaluation on CQ500 and RSNA.. We employ the prompt "This CT study shows: {disease}." during zero-shot transfer.
Figure 3: Results of linear-probe and zero-shot evaluations on the CQ500 [21] and RSNA [22] datasets. We report AUC for each class, with blue represents our HLIP; green represents FM-HeadCT [20]; and red represents Google CT [53] in linear-probe evaluation and ViT in zero-shot evaluation.
Results. Consistent with FM-HeadCT [20], we report the area under the ROC curve (AUC) for each class in Figure 3. In linear-probe evaluation, HLIP demonstrates superior performance over FM-HeadCT [20], which is pre-trained on 361,663 studies using DINOv2 [55], and the Google-CT [53] foundation model. Specifically, on the CQ500 dataset [21], HLIP outperforms both FM-HeadCT and Google-CT with macro AUC improvements of 8.3% and 12.2%, respectively; on the RSNA dataset [22], HLIP achieves macro AUC improvements of 1.7% and 5.8% over the these baselines, respectively. In zero-shot evaluation, HLIP consistently outperforms the vanilla ViT, achieving macro AUC improvements of 10% on CQ500 and 2.4% on RSNA, further demonstrating the effectiveness of the proposed hierarchical attention mechanism.
@@ 4.3 Chest CT
Implementation Details. Following CT-CLIP [3] and fVLM [6], we apply HLIP to the CT-RATE [3] training set, then perform internal validation on its test split and external validation on the full Rad-ChestCT dataset [24]. The implementation details differ from those in Section 3.2. Following prior works [34, 6], we standardize each scan to a spacing of (3mm,1mm,1mm), truncate HU values to [-1150,350], construct mini-batches with a center crop of size (112,336,336), and apply a token size of (8,24,24). We also use CXR-BERT [7] as the text encoder. HLIP is trained for 20 epochs without patch dropout. Using a batch size of 512 on 4 A40 GPUs, training completes within 6 hours, significantly faster than prior works [3, 6], which require several days even with more advanced computational resources. More details are provided in Appendix B.1.
Table 2: Results of multi-label zero-shot evaluation on the CT-RATE [3] test split for internal validation and the Rad-ChestCT [24] for external validation. The best results are highlighted in bold.
@@
CT-CLIP [3] 73.3 66.9 70.8 32.6 - - 63.3 59.9 64.7 34.1 - -
BIUD [34] 71.3 68.1 71.6 33.8 67.3 62.9 60.6 65.2 33.7 59.6
Merlin [4] 72.8 67.2 70.9 33.7 70.1 64.4 61.9 66.3 34.8 61.0
@@
Supervised by Qwen [56] Summarized Reports
fVLM [6] 77.8 71.8 75.1 37.9 72.8 68.0 64.7 68.8 37.4 64.6
@@
Results. Consistent with fVLM [6], we report the area under ROC curve (AUC); balanced accuracy (ACC); weighted F1-score (F1); recall; and precision for multi-label zero-shot classification in Table 2. When pre-trained with original reports, HLIP outperforms second-best models by 4.9% AUC on internal and 7.9% AUC on external validation. When pre-trained with reports summarized by large language models (e.g., Qwen [56], as used by fVLM), HLIP outperforms fVLM by 0.9% and 3.7% AUC on internal and external validation, respectively. These results demonstrate that, compared with other components such as the VQ-VAE [57] tokenizer (used by CT-CLIP) and anatomical-guided fine-grained alignment (used by fVLM), the proposed hierarchical attention is a more effective and generalizable adaptation for language-image pre-training in 3D medical imaging.
@@
Figure 4: Ablation study on the Pub-Brain-5 (a)-(d) and Rad-ChestCT [24] dataset (e)-(h). We report balanced accuracy (ACC) on both datasets.
@@ 4.4 Ablation Study
In Figure 4, we present an ablation study on the key components of our visual encoder, including the cls token propagating strategy (Figure 4(a) and 4(e)); the location of dense attention (Figure 4(b) and 4(f)); and the number of dense attention layers (Figure 4(c) and 4(c)). We additionaly investigate the batch size in Figure 4(d) and 4(h), a factor overlooked in previous work [3, 6].
CLS token. When propagating the cls token across different attention levels, in addition to the default averaging setting, we also consider two alternative scenarios: (i) replacing the cls token with global average pooling at the end of the visual encoder (none); and (ii) applying weighted averaging using an attention pooling (learned). As shown in Figures 4(a) and 4(e), aggregating the cls token by averaging proves to be the most effective approach.
Location of dense attention layers. In Figure 4(b) and 4(f), we consider alternative locations of the dense attention layer: study attention for uncurated datasets, where each study contains multiple scans, and scan attention for curated datasets, where each study contains a single scan. We find that evenly distributed dense attention layers outperform other distributions, such as placing all at the front or all at the back.
Number of dense attention layers. In Figure 4(c) and 4(c), we investigate the number of dense attention layers, where 0 and 12 correspond to two implementations of vanilla ViT. We find that 4 dense attention layers are sufficient to capture global information, outperforming both 0 layers, which lack global information, and 12 layers, which lack the priors provided by slice or scan attention.
Batch size. Previous work [3, 6] used only small batch sizes (e.g., 48) during the pre-training. In Figure 4(d) and 4(h), we investigate this factor, which has been widely recognized as important in language-image pre-training, in the context of 3D medical imaging. We find that language-image pre-training in 3D medical imaging still benefits from larger batch sizes, underscoring the importance of memory-friendly designs such as our hierarchical attention mechanism.
@@ 4.5 Visualizations
Figure 5: Qualitative results of zero-shot diagnosis on the Rad-ChestCT [24] and BraTS [47, 48] datasets. The first row shows the original clinical scans with pathologic regions outlined (dashed red). The second row shows activation HLIP maps [58]. HLIP identifies the pathologic regions across multiple groups of adjacent chest CT slides (left) and brain MRI scans (right).
In Figure 5, we visualize the activation maps [58] of HLIP in the zero-shot setting on Rad-ChestCT [24] and BraTS [47, 48]. For chest CT (curated, single scan), HLIP attends to pathologic regions across different slices, e.g., slice Z and Z+8, by leveraging scan attention. For brain MRI (uncurated, multiple scans), HLIP attends to pathologic regions across different scan types, e.g., FLAIR and T1W+contrast, by leveraging study attention. This demonstrates the effectiveness HLIP in modeling the hierarchical structure of 3D medical imaging.
@@ 4.6 Clinical Translation of HLIP
Figure 6: Prospective testing results on (a) 52 brain MRI diagnoses and (b) 83 head CT diagnoses. Results for each diagnosis are presented in Figures 11 and 12 in Appendix C.
We conduct a 1-year, health system-scale, comprehensive evaluation on a prospective set of neurological studies. The datasets include ˜23K brain MRI studies covering 52 diagnoses and ˜15K head CT studies covering 83 diagnoses. We report macro AUC in Figure 6(a) and 6(b). HLIP consistently outperforms ViT. We provide detailed results in Appendix C.

### v3

@@ 4 Experiments
We apply HLIP to brain MRI, head CT, and chest CT. For brain MRI, we construct a new benchmark for zero-shot classification, using public datasets (Dufumier et al., 2022; Liu et al., 2023; Baid et al., 2021; LaBella et al., 2023; Moawad et al., 2024; Kazerooni et al., 2024; Rudie et al., 2024; Link et al., 2024). For head CT and chest CT, we follow evaluation protocols from previous work (Hamamci et al., 2024; Shui et al., 2025; Zhu et al., 2025). Experiments on our uncurated brain MRI and head CT datasets underscore the importance of effective modeling and demonstrate superior performance of HLIP over current foundation models (Zhang et al., 2023a; Nie et al., 2025; Zhu et al., 2025; Yang et al., 2024). Experiments on curated chest CT datasets (Draelos et al., 2021; Hamamci et al., 2024) further isolate the effectiveness of the proposed hierarchical attention mechanism. Finally, we conduct comprehensive ablation and analysis.
@@ 4.1 Brain MRI
Pub-Brain-5. To the best of our knowledge, no publicly available benchmark exists for zero-shot transfer evaluation on brain MRI tasks [^2] BrainMD (Wang et al., 2024) is not publicly available due to an unexpected privacy policy.. To this end, we construct a benchmark, named Pub-Brain-5, based on existing publicly available brain MRI datasets. Pub-Brain-5 comprises 18,343 studies drawn from Open-BHB (Dufumier et al., 2022), the Stroke dataset (Liu et al., 2023), BraTS 2023 (Baid et al., 2021; LaBella et al., 2023; Moawad et al., 2024; Kazerooni et al., 2024), NYU-Mets (Link et al., 2024), and UCSF-Mets (Rudie et al., 2024). It spans five classes: healthy (3,984), acute stroke (2,871), glioma (1,614), meningioma (1,141), and metastasis (8,733). For each study, the number of scans ranges from 1 to 14. We also construct a subset of Pub-Brain-5, namely Pub-Brain-5-GT, which contains lesion-containing slice annotations derived from the original segmentation ground truth. Pub-Brain-5-GT comprises 8,944 studies covering the same five classes: healthy (3,984), acute stroke (2,372), glioma (1,350), meningioma (1,000), and metastasis (238). We evaluate three zero-shot tasks: (i) binary anomaly detection (i.e., distinguishing pathological from healthy studies); (ii) three-way tumor classification; and (iii) five-way disease classification.
Baselines. We first evaluate the zero-shot transferability of two biomedical CLIP models on our benchmark: BiomedCLIP (Zhang et al., 2023a) and ConceptCLIP (Nie et al., 2025). Since both models require 2D inputs, we generate study predictions by applying average pooling to the outputs across all slices. We acknowledge that slices without lesions introduce noise for these 2D baselines; however, this is an intrinsic limitation of such models. We support these two baselines on Pub-Brain-5-GT, where predictions are made only on lesion-containing slices, while HLIP continues to take the raw study as input. In addition, we evaluate Prima (Lyu et al., 2026), which represents the vqvae + hierarchical transformer architecture family (Hamamci et al., 2024; Lyu et al., 2026), as well as a vanilla ViT that encodes the entire study.
Implementation Details. We use the prompt "This brain MRI shows: {disease}." to perform zero-shot inference (Radford et al., 2021) with biomedical CLIP models. HLIP and the vanilla ViT are trained on BrainMRI220K as described in Section 3.2, while Prima (Lyu et al., 2026) is re-trained on the same dataset using the training recipe provided in the original paper. These three models employ the prompt "This MRI study shows: {disease.}" during zero-shot transfer.
@@
Method Anomaly Detection Tumor Disease
@@
BiomedCLIP (Zhang et al., 2023a) 64.7 87.8 63.6 59.8 69.0 50.4 31.5
ConceptCLIP (Nie et al., 2025) 66.8 91.9 57.7 67.9 71.1 35.7 30.9
Prima (Lyu et al., 2026) 78.8 89.3 70.8 64.7 75.9 42.8 31.6
@@
Results. Given both benchmarks are imbalanced, we report balanced accuracy (ACC) in Table 1. On Pub-Brain-5, HLIP demonstrates superior performance over all baseline models, outperforming the second-best baseline by 20.5% ACC in disease classification. On Pub-Brain-5-GT, we observe a substantial performance boost for two biomedical CLIP models (Zhang et al., 2023a; Nie et al., 2025) when predictions are restricted to lesion-containing slices, underscoring the fairness of our evaluation for these models even on uncurated studies. Notably, in disease classification, ConceptCLIP (Nie et al., 2025) achieves 50.8% ACC, outperforming Prima (Lyu et al., 2026) and the vanilla ViT trained on our large-scale domain-specific datasets, indicating that merely scaling up data is not a sufficient solution for language-image pre-training in 3D medical imaging. On the other hand, HLIP achieves the new SOTA 61.3% ACC, attributable to the effectiveness of the proposed hierarchical attention mechanism.
HLIP can occasionally lag behind BiomedCLIP (Zhang et al., 2023a) and ConceptCLIP (Nie et al., 2025), particularly in the detection of brain glioma. However, we note that datasets like BraTS (Baid et al., 2021) are likely included in their pre-training corpora, as both models are trained on figure-caption pairs from PubMed. HLIP, on the other hand, does not encounter any instances from public datasets during training. The skull-stripping used in BraTS actually exacerbates the distribution shift. Moreover, HLIP makes predictions directly from the original radiology study, offering greater flexibility for real-world applications.
@@ 4.2 Head CT
Implementation Details. We apply HLIP to HeadCT240K, as described in Section 3.2. For linear-probe evaluation, we benchmark HLIP on CQ500 (Chilamkurthy et al., 2018) and RSNA (Flanders et al., 2020), exactly following the train–test split and hyperparameters of FM-HeadCT (Zhu et al., 2025). For zero-shot evaluation, we evaluate HLIP on the full CQ500 and RSNA datasets and compare it with the vanilla ViT also trained on the same HeadCT240K dataset [^3] Recent CLIP models for 3D medical imaging, such as Merlin (Blankemeier et al., 2024) and M3D (Bai et al., 2024), are excluded from our baselines, as they perform no better than random guessing in zero-shot evaluation on CQ500 and RSNA.. We employ the prompt "This CT study shows: {disease}." during zero-shot transfer.
Figure 3: Results of linear-probe and zero-shot evaluations on the CQ500 (Chilamkurthy et al., 2018) and RSNA (Flanders et al., 2020) datasets. We report AUC for each class. In linear-probe, red represents Google CT (Yang et al., 2024); green represents FM-HeadCT (Zhu et al., 2025); and blue represents our HLIP. In zero-shot evaluation, red represents the vanilla ViT and blue represents our HLIP.
Results. Consistent with FM-HeadCT (Zhu et al., 2025), we report the area under the ROC curve (AUC) for each class in Figure 3. In linear-probe evaluation, HLIP demonstrates superior performance over FM-HeadCT (Zhu et al., 2025), which is pre-trained on 361,663 studies using DINOv2 (Oquab et al., 2023), and the Google-CT (Yang et al., 2024) foundation model. Specifically, on the CQ500 dataset (Chilamkurthy et al., 2018), HLIP outperforms both FM-HeadCT and Google-CT with macro AUC improvements of 8.3% and 12.2%, respectively; on the RSNA dataset (Flanders et al., 2020), HLIP achieves macro AUC improvements of 1.7% and 5.8% over the these baselines, respectively. Outperforming current foundation models demonstrates the clinical significance of HLIP. In zero-shot evaluation, HLIP consistently outperforms the vanilla ViT, achieving macro AUC improvements of 10.0% on CQ500 and 2.4% on RSNA, further demonstrating the effectiveness of the proposed hierarchical attention mechanism.
@@ 4.3 Chest CT
Implementation Details. Following CT-CLIP (Hamamci et al., 2024) and fVLM (Shui et al., 2025), we apply HLIP to the CT-RATE (Hamamci et al., 2024) training set, then perform internal validation on its test split and external validation on the full Rad-ChestCT dataset (Draelos et al., 2021). The implementation details differ from those in Section 3.2. Following prior works (Cao et al., 2024; Shui et al., 2025), we standardize each scan to a spacing of (3mm,1mm,1mm), truncate HU values to [-1150,350], construct mini-batches with a center crop of size (112,336,336), and apply a token size of (8,24,24). We also use CXR-BERT (Boecking et al., 2022) as the text encoder. Other than the aforementioned modifications, all remaining settings are exactly the same as those described in Section 3.2. HLIP is trained for 20 epochs without patch dropout. Using a batch size of 512 on 4 A40 GPUs, training completes within 6 hours, significantly faster than prior works (Hamamci et al., 2024; Shui et al., 2025), which require several days even with more advanced computational resources. More details are provided in Appendix B.1.
Table 2: Results of multi-label zero-shot evaluation on the CT-RATE (Hamamci et al., 2024) test split for internal validation and the Rad-ChestCT (Draelos et al., 2021) for external validation. The best results are highlighted in bold.
@@
CT-CLIP (Hamamci et al., 2024) 73.3 66.9 70.8 32.6 - - 63.3 59.9 64.7 34.1 - -
BIUD (Cao et al., 2024) 71.3 68.1 71.6 33.8 67.3 62.9 60.6 65.2 33.7 59.6
Merlin (Blankemeier et al., 2024) 72.8 67.2 70.9 33.7 70.1 64.4 61.9 66.3 34.8 61.0
@@
Supervised by Qwen (Bai et al., 2023) Summarized Reports
fVLM (Shui et al., 2025) 77.8 71.8 75.1 37.9 72.8 68.0 64.7 68.8 37.4 64.6
@@
Results. Consistent with fVLM (Shui et al., 2025), we report the area under ROC curve (AUC); balanced accuracy (ACC); weighted F1-score (F1); recall; and precision for multi-label zero-shot classification in Table 2. When pre-trained with original reports, HLIP outperforms second-best models by 4.9% AUC on internal and 7.9% AUC on external validation. When pre-trained with reports summarized by large language models (e.g., Qwen (Bai et al., 2023), as used by fVLM), HLIP outperforms fVLM by 0.9% and 3.7% AUC on internal and external validation, respectively. These results demonstrate that, compared with other components such as the VQ-VAE (Van Den Oord et al., 2017) tokenizer (used by CT-CLIP) and anatomical-guided fine-grained alignment (used by fVLM), the proposed hierarchical attention is a more effective and generalizable adaptation for language-image pre-training in 3D medical imaging.
@@
Figure 4: Ablation study on the Pub-Brain-5 (a)-(d) and Rad-ChestCT (Draelos et al., 2021) dataset (e)-(h). We report balanced accuracy (ACC) on both datasets.
@@ 4.4 Ablation Study
In Figure 4, we present an ablation study on the key components of our visual encoder, including the cls token propagating strategy (Figure 4(a) and 4(e)); the location of dense attention (Figure 4(b) and 4(f)); and the number of dense attention layers (Figure 4(c) and 4(c)). We additionaly investigate the batch size in Figure 4(d) and 4(h), a factor overlooked in previous work (Hamamci et al., 2024; Shui et al., 2025).
CLS token. When propagating the cls token across different attention levels, in addition to the default averaging setting, we also consider two alternative scenarios: (i) global average pooling at the end of the visual encoder (none); and (ii) applying weighted averaging using an attention pooling (learned). As shown in Figures 4(a) and 4(e), aggregating the cls token by averaging proves to be the most effective approach. This result demonstrates that more complex components are not necessary to achieve better performance.
Location of dense attention layers. In Figure 4(b) and 4(f), we consider alternative locations of the global attention layer (i.e., study attention for uncurated datasets and scan attention for curated datasets). Compared with distributions such as placing all global attention layers at the front or all at the end, we find that evenly distributed global attention layers yield the best performance. This is because they enable the model to capture global information at multiple feature levels.
Number of dense attention layers. In Figure 4(c) and 4(c), we investigate the number of global attention layers, where 0 and 12 correspond to two implementations of vanilla ViT. We find that 4 global attention layers are sufficient to capture global information, outperforming both 0 layers, which lack global information, and 12 layers, which lack the constructive priors provided by slice or scan attention.
Batch size. Previous work (Hamamci et al., 2024; Shui et al., 2025) used only small batch sizes (e.g., 48) during the pre-training. In Figure 4(d) and 4(h), we investigate this factor, which has been widely recognized as important in language-image pre-training, in the context of 3D medical imaging. We find that language-image pre-training in 3D medical imaging still benefits from larger batch sizes, underscoring the importance of memory-friendly designs such as our hierarchical attention mechanism.
@@ 4.5 Visualizations
Figure 5: Qualitative results of zero-shot diagnosis on the Rad-ChestCT (Draelos et al., 2021) and BraTS (Baid et al., 2021; LaBella et al., 2023) datasets. The first row shows the original clinical scans with pathologic regions outlined (dashed red). The second row shows activation maps (Chefer et al., 2021). HLIP identifies the pathologic regions across multiple groups of adjacent chest CT slides (left) and brain MRI scans (right).
In Figure 5, we visualize the activation maps of HLIP, following  Chefer et al. (2021) in the zero-shot setting on Rad-ChestCT (Draelos et al., 2021) and BraTS (Baid et al., 2021; LaBella et al., 2023). Although HLIP mainly employs lightweight local attention layers, these visualization results demonstrate the effectiveness of a few global attention layers in enabling HLIP to attend to visual features globally. Specifically, for chest CT (curated, single scan), HLIP attends to pathologic regions across different slices, e.g., slice Z and Z+8, by leveraging scan attention. For brain MRI (uncurated, multiple scans), HLIP attends to pathologic regions across different scan types, e.g., FLAIR and T1W+contrast, by leveraging study attention. This demonstrates the effectiveness HLIP in modeling the hierarchical structure of 3D medical imaging.
@@ 4.6 Clinical Translation of HLIP
Figure 6: Prospective testing results on (a) 52 brain MRI diagnoses and (b) 83 head CT diagnoses.
We conduct a 1-year, health system-scale, comprehensive evaluation on a prospective set of neurological studies. The datasets include ˜23K brain MRI studies covering 52 diagnoses and ˜15K head CT studies covering 83 diagnoses. We report macro AUC in Figure 6(a) and 6(b). HLIP consistently outperforms ViT. Results for each diagnosis are presented in Figures 11 and 12 in Appendix C.